### v1

@@ 5.1 PLD and $f$-DP filters are NOT free in general
Consider two tradeoff curves that cross $g_{1}=f_{\varepsilon_{1},\delta_{1}}\in\mathcal{F}$ and $g_{2}=f_{\varepsilon_{2},\delta_{2}}\mathcal{F}$, that fall in case 1 in the proof of Proposition˜13 (this is possible, but the proof is identical for case 2, swapping the proper slope).
Pick $f=f_{\varepsilon,\delta}$ with $\varepsilon=\log\big(e^{2\varepsilon_{1}}-(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)$ (that is, $e^{-\varepsilon}$ is the first slope of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$) and $\delta=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$ (that is, the $\delta$ of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$).
Consider an adversary playing Algorithm˜2 with a $g_{1}$ mechanism first, and then an adaptive choice between $g_{1},g_{2}$. By construction, $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})\geq f$, so $g_{1}\otimes g_{1}\geq f$ and $g_{1}\otimes g_{2}\geq f$: Algorithm˜2 always accepts the query.
From the proof of Theorem˜2 we know that a two-steps filter can play $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$. By construction, we have that $f(\alpha)=\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)$ on $\alpha\in(0,\alpha_{1}^{\star 2}]$ From Figure˜1(a) (and the proof of Proposition˜13) we know that on $\alpha\in(0,\alpha_{1}^{\star 2}]$, we have $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)=f(\alpha)$. Algorithm˜2 is thus not $f$-DP, concluding the proof. ∎
(a) GDP $G_{\mu}$ filter bound
(b) Approximate GDP $f_{0,\delta}\otimes G_{\mu}$ filter bound
Figure 3: Counter example to free $f$-DP filters when $f$ is GDP (a) or approximate GDP (b).
Corollary 3.
The set of all tradeoff curves $\mathcal{T}$ does not have free $f$-DP filters even for $f\in\{G_{\mu}:\mu\geq 0\}$.
The set of $(\varepsilon,\delta)$-DP tradeoff curves $\mathcal{F}=\{f_{\varepsilon,\delta}:\varepsilon\geq 0,\ \delta\in[0,1]\}$ does not have free $f$-DP filters even for $f\in\{f_{0,\delta}\otimes G_{\mu}:\mu\geq 0,\delta\in[0,1]\}$.
@@
Following the reasoning of Corollary˜2, we show that we can construct a set of two tradeoff curves $g_{1},g_{2}\in\mathcal{F}$, and a bound $f$ in the correct family, such that $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})\geq f$ but there exists values of $\alpha$ s.t. $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<f$. Figure˜3(a) shows such a construction for $f\in\{G_{\mu}:\mu\geq 0\}$ and Figure˜3(b) for $f\in\{f_{0,\delta}\otimes G_{\mu}:\mu\geq 0,\delta\in[0,1]\}$. All trade-off curves are (compositions of) piecewise linear functions, or Gaussian trade-offs. They can thus be computed exactly in closed-form. ∎
5.2 Families of mechanisms with free filters
Theorem˜3, combined with Proposition˜10, show that a family of DP mechanisms that is closed under composition, and with well ordered tradeoff curves, admits free $f$-DP (and PLD by Proposition˜10) filters.
Corollary 4.
The set of all Gaussian tradeoff curves $G=\{G_{\mu}:\mu\geq 0\}$ has free filters. That is Algorithm˜2 is $f$-DP.
@@ 5.2 Families of mechanisms with free filters
For any $\mu_{1},\mu_{2}\geq 0$, $G_{\mu_{1}}\otimes G_{\mu_{2}}=G_{\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}}\in G$. If $\mu_{1}>\mu_{2}$, then $G_{\mu_{1}}\preceq G_{\mu_{2}}$. Gaussian mechanisms are closed under composition and well ordered: by Theorem˜3, they have free $f$-DP/PLD filters, as already shown in [koskelaindividual, smith2022fullyadaptivecompositiongaussian]. ∎
Corollary 5.
The set of all $0,\delta$-DP tradeoff curves $\mathcal{H}=\{f_{0,\delta}:\delta\in[0,1]\}$ has free filters. That is Algorithm˜2 if $f$-DP.
@@
For any $\delta_{1},\delta_{2}\geq 0$, $f_{0,\delta_{1}}\otimes f_{0,\delta_{2}}=f_{0,\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}}\in\mathcal{H}$. If $\delta_{1}>\delta_{2}$, then $f_{0,\delta_{1}}\preceq f_{0,\delta_{2}}$. Pure $\delta$ mechanisms are closed under composition and well ordered: by Theorem˜3, they have free $f$-DP/PLD filters. ∎
6 A Proof of Theorem˜3

### v2

@@ 7.3.1 Useful Lemmata
We have
$H_{x}(P\parallel Q)$ $=\sup_{E\subseteq\Omega}P(E)-x\cdot Q(E)$
$=\sup_{E\subseteq\Omega}1-P(\Omega\setminus E)-x\cdot Q(E)$
$\leq\sup_{E\subseteq\Omega}1-P^{\prime}(\Omega\setminus E)-x\cdot Q^{\prime}(E)$
$=\sup_{E\subseteq\Omega}P^{\prime}(\omega^{*})+P^{\prime}(E)-x\cdot Q^{\prime}(E)$
$\leq P^{\prime}(\omega^{*})+H_{x}(P^{\prime}\parallel Q^{\prime})$
$\leq\gamma+H_{x}(P^{\prime}\parallel Q^{\prime}).\qed$
The next lemma asserts that, to check if a symmetric PLD dominates another, it suffices to check the hockey-divergence of order $x\geq 1$. This is unlike the (asymmetric) case where we need to check $x\in(0,1)$ as well. This will be convenient for our proofs. We defer its (simple) proof to Appendix˜C.
Lemma 8.
For any symmetric PLDs $^{\prime}, LLL\succeq L^{\prime}$ iff $h_{x}(L)\geq h_{x}(L^{\prime})$ for all $x\geq 1$.
We will use the following “Multiplicative Azuma’s Inequality” from (Kuszmaul and Qi, 2021, Corollary 11). This is an extension of the classic multiplicative Chernoff bound, with the difference that the distributions $\mathcal{D}_{i}$ can depend on the previous outcomes $W_{1},\dots,W_{i-1}$. The only constraint required is that $\sum_{i\in[k]}\mathop{\mathbb{E}}_{W_{i}\sim\mathcal{D}_{i}}[W_{i}]\leq\mu$ where $\mu$ is a prespecified parameter.
Theorem 16 (Multiplicative Azuma’s Inequality Kuszmaul and Qi (2021)).
Let $\mu>0$. Suppose that an adversary constructs random variables $W_{1},\dots,W_{k}\in[0,1]$ via the following process: When the outcomes of $W_{1},\dots,W_{i-1}$ are determined, the adversary can select the distribution $\mathcal{D}_{i}$ with mean $\mu_{i}$ for $W_{i}$ under the constraint $\sum_{i\in[k]}\mu_{i}\leq\mu$.
If $W=\sum_{i}W_{i}$, then for any $\zeta>0$,
$\Pr[W\geq(1+\zeta)\mu]\leq\exp\left(-\frac{\zeta^{2}}{2+\zeta}\mu\right).$
It will also be convenient to state the following version of the tail bound for a similar setting as above but when $\mu$ is very small.
Lemma 9.
Let the setting be exactly the same as in Theorem˜16. Then, $\Pr[W\geq\mu/\zeta]\leq\zeta$.
@@
This is simply an application of Markov’s inequality, since $\mathop{\mathbb{E}}[W]\leq\mu$. ∎
7.3.2 Privacy Analysis of RR: Comparing Different $\varepsilon$’s
Before we proceed to the main argument, we need some technical results. The main lemma below states that, if we take $\tilde{\varepsilon}-$DP randomized response and compose it 12 times, then its pair dominates that of $2\tilde{\varepsilon}-$DP randomized response. This result seems interesting on its own and might be useful beyond the context of this work. We remark that we did not attempt to optimize the constant “12” here; nevertheless, it is not hard to see that any constant less than 6 does not work (by taking $\varepsilon\to 0$ and looking at the hockey-stick divergence of order $x=1$, i.e. the TV distance).
Lemma 10.
For any $\tilde{\varepsilon}\geq 0$, $(\mathrm{RR}^{\otimes 12}_{\tilde{\varepsilon}}(0),\mathrm{RR}^{\otimes 12}_{\tilde{\varepsilon}}(1))\succeq(\mathrm{RR}_{2\tilde{\varepsilon}}(0),\mathrm{RR}_{2\tilde{\varepsilon}}(1))$.
@@ 7.3.2 Privacy Analysis of RR: Comparing Different $\varepsilon$’s
Let $t=12$. Since both $(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0),\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$ and $(\mathrm{RR}_{2\tilde{\varepsilon}}(0),\mathrm{RR}_{2\tilde{\varepsilon}}(1))$ are symmetric, by Lemma˜8, it suffices to show that $H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))\leq H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$ for any $x\geq 1$. To see this is true, consider three cases:
  • Case I: $x\geq e^{2\tilde{\varepsilon}}$. Then, $H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))=0\leq H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$.
  • Case II: $x=1$. Note that $H_{x}$ is simply the TV distance. Thus, $H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))=\frac{e^{2\tilde{\varepsilon}}-1}{e^{2\tilde{\varepsilon}}+1}.$
    To lower bound $H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$, let $S_{i}$ denote the set of all tuples in $\{0,1\}^{t}$ such that the number of ones is exactly $i$. Let $P=\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0),Q=\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1)$ We thus have
    $H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$ $\geq\sum_{i=0}^{t/2-1}{\left({P(S_{i})-Q(S_{i})}\right)}$
    $=\sum_{i=0}^{t/2-1}\binom{t}{i}\left(\frac{e^{(t-i)\tilde{\varepsilon}}-e^{i\tilde{\varepsilon}}}{(e^{\tilde{\varepsilon}}+1)^{t}}\right)$
    $=\sum_{i=0}^{t/2-1}\binom{t}{i}\left(\frac{e^{i\tilde{\varepsilon}}{\left({e^{(t-2i)\tilde{\varepsilon}}-1}\right)}}{(e^{\tilde{\varepsilon}}+1)^{t}}\right)$
    $\geq\frac{e^{2\tilde{\varepsilon}}-1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({\binom{t}{t/2-1}e^{(t/2-1)\tilde{\varepsilon}}+\sum_{i=0}^{t/2-2}\binom{t}{i}{\left({e^{i\tilde{\varepsilon}}{\left({e^{(t-2i-2)\tilde{\varepsilon}}+1}\right)}}\right)}}\right)}.$
    Thus, we have
    $\frac{H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))}{H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))}$
    $=\frac{e^{2\tilde{\varepsilon}}+1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({\binom{t}{t/2-1}e^{(t/2-1)\tilde{\varepsilon}}+\sum_{i=0}^{t/2-2}\binom{t}{i}{\left({e^{i\tilde{\varepsilon}}{\left({e^{(t-2i-2)\tilde{\varepsilon}}+1}\right)}}\right)}}\right)}$
    $=\frac{1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({\binom{t}{t/2-1}{\left({e^{(t/2+1)\tilde{\varepsilon}}+e^{(t/2-1)\tilde{\varepsilon}}}\right)}+\sum_{i=0}^{t/2-2}\binom{t}{i}{\left({e^{(t-i)\tilde{\varepsilon}}+e^{i\tilde{\varepsilon}}+e^{(t-i-2)\tilde{\varepsilon}}+e^{(i+2)\tilde{\varepsilon}}}\right)}}\right)}$
    $\geq\frac{1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({\binom{t}{t/2-1}{\left({e^{(t/2+1)\tilde{\varepsilon}}+e^{(t/2-1)\tilde{\varepsilon}}}\right)}+\sum_{i=0}^{t/2-2}\binom{t}{i}{\left({e^{(t-i)\tilde{\varepsilon}}+e^{i\tilde{\varepsilon}}}\right)}+2\binom{t}{t/2-2}e^{(t/2)\tilde{\varepsilon}}}\right)}$
    $=\frac{1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({(e^{\tilde{\varepsilon}}+1)^{t}+{\left({2\binom{t}{t/2-2}-\binom{t}{t/2}}\right)}e^{(t/2)\tilde{\varepsilon}}}\right)}$
    $\geq 1,$
    where the last inequality is due to our choice $t=12$.
  • Case III: $x\in{\left({1,e^{2\tilde{\varepsilon}}}\right)}$. Since the privacy loss values for both $\operatorname{PLD}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0),\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$ and $\operatorname{PLD}(\mathrm{RR}_{2\tilde{\varepsilon}}(0),\mathrm{RR}_{2\tilde{\varepsilon}}(1))$ are multiples [^6] Note that for $\operatorname{PLD}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0),\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$, this follows from the fact that $t$ is even. of $2\tilde{\varepsilon}$, the hockey stick curves $H_{x}$ of both pairs are linear for $x\in\left[1,e^{2\tilde{\varepsilon}}\right$]. Thus, we have
    $H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$
    $=\frac{x-1}{e^{2\tilde{\varepsilon}}-1}\cdot H_{e^{2\tilde{\varepsilon}}}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))+\frac{e^{2\tilde{\varepsilon}}-x}{e^{2\tilde{\varepsilon}}-1}\cdot H_{1}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$
    $\geq\frac{x-1}{e^{2\tilde{\varepsilon}}-1}\cdot H_{e^{2\tilde{\varepsilon}}}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))+\frac{e^{2\tilde{\varepsilon}}-x}{e^{2\tilde{\varepsilon}}-1}\cdot H_{1}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))$
    $=H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1)),$
    where the inequality follows from Cases I and II above. ∎
Using the above lemma together with our setting of parameters, it is simple to derive the following, which will be more convenient in a subsequent usage.
Lemma 11.
For every $i\in[n]$, $\mathrm{RR}_{\varepsilon_{i}}^{\otimes b^{*}_{i}}$ is $\left(2\lceil\log(1/\delta)\rceil\cdot(9\varepsilon+2),32\lceil\log(1/\delta)\rceil\cdot\delta\right)$-DP.
@@
Consider three cases:
  • Case I: $i=n$. This case is obvious since $b^{*}_{i}=0$.
  • Case II: $i=1$. In this case, $\ell^{*}_{i}=k$ and $\varepsilon_{i}\leq\frac{\varepsilon}{k}$. Thus, by our setting of $b^{*}_{i}=2\lceil\log(1/\delta)\rceil\cdot\ell^{*}_{i}$ and basic composition, $\mathrm{RR}^{\otimes b_{i}}_{\varepsilon_{i}}$ is $\left({2\lceil\log(1/\delta)\rceil\cdot\varepsilon}\right)$}-DP as desired.
  • Case III: $1<i<n$. From the setting of $\ell^{*}_{i}$ in (1), $\mathrm{RR}_{\varepsilon_{i}}^{\otimes\frac{\ell^{*}_{i}-1}{8}}$ is $(\varepsilon,2\delta)$-DP. Since $\varepsilon_{i}\leq\varepsilon+2$, by basic composition, $\mathrm{RR}_{\varepsilon_{i}}^{\otimes\ell^{*}_{i}}$ is $(9\varepsilon+2,16\delta)$-DP. Thus, by our setting of $b^{*}_{i}=2\lceil\log(1/\delta)\rceil\cdot\ell^{*}_{i}$ and basic composition, $\mathrm{RR}^{\otimes b_{i}}_{\varepsilon_{i}}$ is $\left(2\lceil\log(1/\delta)\rceil\cdot(9\varepsilon+2),32\lceil\log(1/\delta)\rceil\cdot\delta\right)$-DP as desired. ∎
7.3.3 Formulation as Post-Processing of RR
As alluded to earlier, we will prove Theorem˜15 by showing that, with high probability, the filter’s output is a post-processing of randomized response answers. To do this, recall from Section˜2 that it suffices to consider two neighbouring datasets $D_{0},D_{1}$. Our post-processing procedure is described in full in Algorithm˜4. It takes as input the randomized response outputs on a bit $h\in\{0,1\}$, and its output (approximately) matches the output of the additive filter. To formalize this, let $M^{\mathrm{discrete}}_{\mathcal{A}}(D)$ denote the output distribution of the Additive Discretized Privacy Loss Filter (Algorithm˜3) on input dataset $D$ with analyst $\mathcal{A}$ (where other parameters are omitted from the notation for readability).
Algorithm 4 Post-processing Algorithm
0: Discretization points $\bm{\varepsilon}\in\mathcal{E}_{n}$, budgets $\bm{\ell}\in\mathbb{R}^{n}_{\geq 0}$, analyst $\mathcal{A}$, query capacity $k$, neighbouring datasets $D_{0},D_{1}$, independent samples $y_{i}^{j}\sim\mathrm{RR}_{\varepsilon_{i}}(h)$ (where $h\in\{0,1\}$)
$\mathbf{j}\leftarrow\mathbf{0}$ $\triangleright$ $n$-dimensional all-zero vector; indices for current $R$ samples
for $t\leftarrow 1,\dots,k$ do
  $\mathcal{A}(Y_{1},\dots,Y_{t-1})$ gives mechanism $M_{t}$ that has PLD $L_{t}\in\mathcal{L}_{\mathrm{sym}}$
  $\mathbf{p}^{t}\leftarrow$ Discretized privacy loss profile of $L_{t}$ w.r.t. $\bm{\varepsilon}$
  $\bm{\ell}\leftarrow\bm{\ell}-\mathbf{p}^{t}$
  $\kappa_{t}\leftarrow$ Markov kernel with $\kappa_{t}{\left({\mathrm{RR}_{\mathbf{p}^{t},\bm{\varepsilon}}(h^{\prime})}\right)}=M_{t}(D_{h^{\prime}})$ for $h^{\prime}\in\{0,1\}$ $\triangleright$ exists due to Theorem˜11
  if $\bm{\ell}\geq\mathbf{0}$ then
   $i_{t}\leftarrow$ sample according to $\mathbf{p}^{t}$
   $j_{i_{t}}\leftarrow j_{i_{t}}+1$
   if $j_{i_{t}}>b_{i_{t}}$ then
    return FAIL $\triangleright$ Exceed budget; fail and terminate the algorithm
   Sample $Y_{t}\sim\kappa_{t}{\left({i_{t},y^{j_{i_{t}}}}\right)}$
  else
   $Y_{t}\leftarrow\bot$
return $(Y_{1},\ldots,Y_{k})$
Our post-processing procedure (Algorithm˜4) has the parameter $\mathbf{b}=(b_{1},\dots,b_{n})$ representing the “sample budgets” (i.e. how many randomized response samples are allowed to be used). To analyze the algorithm, first note that when there is no sample budgets (i.e. $b_{1}=\cdots=b_{n}=\infty$), the output is exactly the same as the Additive Discretized Privacy Loss Filter:
Observation 17.
When we set $b_{i}=\infty$ for all $i\in[n]$, the output of Algorithm˜4 has exactly the same distribution as $M^{\mathrm{discrete}}_{\mathcal{A}}(D_{h})$.
Proof.
Since $b_{i}=\infty$ for all $i\in[n]$, we never output FAIL. Thus, for each $t\in[k]$, $(i_{t},y^{j_{i_{t}}})$ has exactly the same distribution as $\mathrm{RR}_{\mathbf{p}^{t},\bm{\varepsilon}}(h)$. From the definition of $\kappa_{t}$, $Y_{t}\sim\kappa_{t}(i_{t},y^{j_{i_{t}}})$ has exactly the same distribution as $M_{t}(D_{h})$. Thus, the entire process is exactly the same as $M^{\mathrm{discrete}}_{\mathcal{A}}(D_{h})$. ∎
We next show that, even when we set the budget as in (1), the algorithm barely fails:
Lemma 12.
When we set $\ell_{i}=\ell^{*}_{i}$ and $b_{i}=b^{*}_{i}$ for all $i\in[n]$ (where $\bm{\ell}^{*},\mathbf{b}^{*}$ are as specified in (1) and (5), respectively), Algorithm˜4 outputs FAIL with probability at most $4n\cdot\delta$.
Proof.
The algorithm outputs FAIL only when $j_{i}>b_{i}$ for some $i\in[n]$. Thus, by the union bound, it suffices to show that $\Pr[j_{i}\geq b_{i}]\leq 4\delta$ for every $i\in[n]$. To do this, we consider two cases:
  • Case I: $i<n$. Let $\mathcal{D}_{t}$ denote the distribution $\mathrm{Ber}(p^{t}_{i})$ and $W_{t}$ denote $\mathbf{1}[i_{t}=i]$. For notational convenience, we also assume that, if the algorithm outputs FAIL in the $t_{0}$-th iteration, then we simply set $p^{t}_{i}=0,\mathcal{D}_{t}=\mathrm{Ber}(0)$ for $t\in\{t_{0}+1,\dots,k\}$.
    Under these notations, we have $W_{t}\sim\mathcal{D}_{t}$ and $W=W_{1}+\cdots+W_{k}$ is exactly $j_{i}$. Finally, due to the constraint of the filter, we have $\mathop{\mathbb{E}}[W_{1}]+\cdots+\mathop{\mathbb{E}}[W_{k}]=p^{1}_{i}+\cdots+p^{k}_{i}\leq\ell^{*}_{i}$. Thus, applying Theorem˜16 with $\zeta=6\lceil\log(1/\delta)\rceil-1$ and observing that $\frac{\zeta^{2}}{2+\zeta}\geq\frac{\zeta+1}{6}\geq\lceil\log(1/\delta)\rceil$, we have
    $\Pr\left[j_{i}>b^{*}_{i}\right]=\Pr[W>(1+\zeta)\cdot\ell^{*}_{i}]\leq\exp(-\log(1/\delta))=\delta.$
  • Case II: $i=n$. Using the same notation as before and recalling that $b_{i}^{*}=0$ and $\ell^{*}_{i}=4\delta$ in this case, we have
    $\Pr\left[j_{i}>b^{*}_{i}\right]=\Pr[j_{i}\geq 1]=\Pr[W\geq 1]$ $\leq\ell^{*}_{i}=4\delta,$
    where the inequality is due to Lemma˜9. ∎
7.3.4 Putting Things Together: Proof of Theorem˜15
Combing all the pieces together, we can now prove Theorem˜15.
Proof of Theorem˜15.
Let $\varepsilon^{\prime}=n\cdot 2\lceil\log(1/\delta)\rceil\cdot(9\varepsilon+2)$ and $\delta^{\prime}=n\cdot 32\lceil\log(1/\delta)\rceil\cdot\delta+4n\delta$. We will show that the $\operatorname{PLD}{\left({\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}^{*}}^{k}}\right)}\preceq R_{\varepsilon^{\prime},\delta^{\prime}}$.
Let $D_{0},D_{1}$ be any neighbouring input datasets. From Observation˜17, $M^{\mathrm{discrete}}_{\mathcal{A}}(D_{h})$ is exactly the distribution of the output of Algorithm˜4 with $b_{i}=\infty$. Consider also the output distribution of Algorithm˜4 with $b_{i}=b^{*}_{i}$ as defined in (5); let us call the resulting distribution $P^{\prime}(D_{h})$.
By Lemma˜12, $(P,Q)=(M^{\mathrm{discrete}}_{\mathcal{A}}(D_{0}),M^{\mathrm{discrete}}_{\mathcal{A}}(D_{1}))$ and $(P^{\prime},Q^{\prime})=(P^{\prime}(D_{0}),P^{\prime}(D_{1}))$ satisfies the conditions in Lemma˜7 with $\omega^{*}=\text{FAIL}$ and $\gamma=4n\delta$. Thus, Lemma˜7 implies that
$H_{e^{\varepsilon^{\prime}}}(M^{\mathrm{discrete}}_{\mathcal{A}}(D_{0})\parallel M^{\mathrm{discrete}}_{\mathcal{A}}(D_{1}))\leq H_{e^{\varepsilon^{\prime}}}(P^{\prime}(D_{0})\parallel P^{\prime}(D_{1}))+4n\delta.$
To bound $H_{e^{\varepsilon^{\prime}}}(P^{\prime}(D_{0})\parallel P^{\prime}(D_{1}))$, notice that, since we set $b_{i}=b^{*}_{i}$ for all $i\in[n]$, $P^{\prime}(D_{h})$ is simply a post-processing of $\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(h)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(h)$. Thus, we have
$H_{e^{\varepsilon^{\prime}}}(P^{\prime}(D_{0})\parallel P^{\prime}(D_{1}))$ $\leq H_{e^{\varepsilon^{\prime}}}\left(\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(0)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(0)~\middle\|~\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(1)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(1)\right).$
Finally, from Lemma˜11 and basic composition, we can conclude that $\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}$ is $\left({\varepsilon^{\prime},n\cdot 32\lceil\log(1/\delta)\rceil\cdot\delta}\right)$}-DP. In other words,
$H_{e^{\varepsilon^{\prime}}}\left(\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(0)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(0)~\middle\|~\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(1)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(1)\right)\leq n\cdot 32\lceil\log(1/\delta)\rceil\cdot\delta.$
Combining the three inequalities, we get
$H_{e^{\varepsilon^{\prime}}}(M^{\mathrm{discrete}}_{\mathcal{A}}(D_{0})\parallel M^{\mathrm{discrete}}_{\mathcal{A}}(D_{1}))\leq n\cdot 32\lceil\log(1/\delta)\rceil\cdot\delta+4n\delta=\delta^{\prime}.$
Thus, $\operatorname{PLD}{\left({\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}^{*}}^{k}}\right)}\preceq R_{\varepsilon^{\prime},\delta^{\prime}}$ as claimed. ∎
8 A Proof of Lemma˜4